OpenStudy (anonymous):

how do youu simplify the radical 96

OpenStudy (zehanz):

96 = 16 * 6, so you could wite it as\[\sqrt{16*6}=\sqrt{16}*\sqrt{6}\] Now there is only one step left for you to do...